OpenStudy (anonymous):

Use Pythagorean identity cos^2(theta) + sine^2(thetaof theta cosine

OpenStudy (anonymous):

Well, you can certainly isolate \(\cos\theta\), but you need more than just that identity to find it.

OpenStudy (anonymous):

sin (theta) = sqrt(7)/7

OpenStudy (anonymous):

\[ \sin^2\theta + \cos^2\theta = 1\implies |\cos\theta| = \sqrt{1-\sin^2\theta} \]Since we know that \(\theta\) is acute, we know that \(\cos\theta > 0\), and so \(|\cos\theta| = \cos\theta\).

OpenStudy (anonymous):

Given the equation you just gave, we have: \[ \cos\theta = \sqrt{1- \left(\frac{\sqrt 7}{7}\right)^2} \]

OpenStudy (anonymous):

You just need to simplify.

OpenStudy (anonymous):

I'm having trouble with the simplification, sorry.

OpenStudy (anonymous):

Start simple: \[ \left(\frac{\sqrt 7}{7}\right)^2=\left(\frac{\sqrt 7}{7}\right)\left(\frac{\sqrt 7}{7}\right)=\frac{\sqrt 7\cdot \sqrt 7}{7\cdot 7} =\ldots \]Can you simplify this?

OpenStudy (anonymous):

I'm getting sqrt(6)/7

OpenStudy (anonymous):

Not quite it

OpenStudy (anonymous):

the answer to your last question is 1/7

OpenStudy (anonymous):

What is \(1 - 1/7\)?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

No, \[ 1-\frac17 \]

OpenStudy (anonymous):

\[ 1-\frac 17 = \frac77-\frac17 = \ldots \]

OpenStudy (anonymous):

I'm getting 6/7

OpenStudy (anonymous):

Okay, so: \[ \sqrt{\frac 67} = \frac{\sqrt 6}{\sqrt7} \]To rationalize the denominator, you multiply top and bottom by it: \[ \frac{\sqrt 6}{\sqrt7}= \frac{\sqrt 6\cdot\sqrt7}{\sqrt7\cdot\sqrt7} = \ldots \]

OpenStudy (anonymous):

sqrt(42)/7

OpenStudy (anonymous):

That looks right.

OpenStudy (anonymous):

\[ \left(\frac{\sqrt{7}}7\right)^2+\left(\frac{\sqrt{42}}7\right)^2 = 1 \]

OpenStudy (anonymous):

Also, \(\sqrt{42}/2 < 1 < \pi /2\), so it is acute.

OpenStudy (anonymous):

thank you